OpenStudy (samigupta8):

Let A and B be two nonsingular square matrices of same order such that B is mot an identity matrix and AB^2=BA.If A^3=B^-1A^3B^n then value of n is

OpenStudy (samigupta8):

B^-1 here means the inverse of B

OpenStudy (samigupta8):

@ganeshie8

OpenStudy (knov):

\[A ^{3}=B ^{-1}*A ^{3}*B ^{n}\] is what you have written yes ?

OpenStudy (samigupta8):

Yeah!

OpenStudy (knov):

if AB^2=BA, then A^3=B^-1A^3B^n which is n here equals 1

OpenStudy (samigupta8):

Where did u use the first relation Ab^2=BA?

OpenStudy (samigupta8):

Options are 1)4 2).2 3).8 4).9

OpenStudy (samigupta8):

@kainui

OpenStudy (kainui):

I left multiplied by the inverse of B to obtain: \[B^{-1}AB^2 = A\] then I cubed this. Then I used \(AB^2=BA\) to squeeze the A's together since this basically lets you commute A with B. Play around with it, give it your best shot. @knov already gave you a great answer.

OpenStudy (samigupta8):

Bt we are not having an option as such for n=1..

OpenStudy (samigupta8):

I m getting n=2.. And it's not right ..why so ??

OpenStudy (samigupta8):

@kainui i even got n=4 .. Are there many values of n for which it can be true??

OpenStudy (kainui):

I don't think so since A and B are nonsingular. Can you show your work for how you got it to get to two different things? Maybe there's a mistake in your work I can help you find.

OpenStudy (samigupta8):

Okay !! And yes it was a mistake on my part .. I just got n=2..

OpenStudy (samigupta8):

Shall i show how i got this ??

OpenStudy (samigupta8):

@kainui

OpenStudy (kainui):

Yeah sure if you like I will happily check your work :D

OpenStudy (samigupta8):

A^3=B^-1A^3B^n Multiplying by B I get, BA^3=A^3B^n (BA)A^2=A^3B^n AB^2A^2=A^3B^n A^3B^2=A^3B^n

OpenStudy (samigupta8):

@kainui

OpenStudy (kainui):

I followed all your steps but I don't see how you did the very last step: AB^2A^2=A^3B^n A^3B^2=A^3B^n Are you saying that \(A^2B^2=B^2A^2\) ?

OpenStudy (samigupta8):

No i applied associativity ..like (AB)C=A(BC)

OpenStudy (samigupta8):

Okk ..we don't have to do like that since what u told is definitely right.. Bt how to solve next??

OpenStudy (kainui):

Since BA=AB^2 you can slowly push these Bs through A one at a time: AB^2A^2=A^3B^n Just rewriting to be clearer before I substitute: AB(BA)A=A^3B^n Now substituting: ABAB^2A=A^3B^n now we do the same with this here: A(BA)B^2A=A^3B^n AAB^2B^2A=A^3B^n A^2B^4A=A^3B^n So now we have 2 of the As over on the left, now you just gotta push that farthest right A through the 4 Bs to the left of it by this same process.

OpenStudy (samigupta8):

Okay ! Pls wait for a sec or so!! I m checking..

OpenStudy (samigupta8):

I got this till here!! Now i m making something more substitution further...

OpenStudy (samigupta8):

I got the ans 8

OpenStudy (samigupta8):

Thanks @kainui

OpenStudy (kainui):

Yeah that sounds right to me :D